# Exceptional Points in the Scattering Resonances of a Sphere Dimer

### v1

@@ IV.1 Exceptional Points
Figure 3: (a) Normalized eigenfrequency deviation ${\delta\omega_{\mathrm{qs}}}/{\omega_{\mathrm{ep}}^{\mathrm{qs}}}={(\omega_{\pm}-\omega_{\mathrm{ep}}^{\mathrm{qs}})}/{\omega_{\mathrm{ep}}^{\mathrm{qs}}}$ in the complex-frequency plane for varying $\gamma_{2}=\gamma=-\gamma_{1}$, comparing full-wave (FW, blue) and quasistatic (QS, red) models for a conjugate-symmetric ($\chi_{1}=\chi_{2}^{*}$) dimer of Drude spheres with $k_{\mathrm{p}1}a=k_{\mathrm{p}2}a=0.1$ and $d=5a$. (b) Normalized eigenfrequency deviation in the FW regime ${\delta\omega}/{\omega_{\mathrm{ep}}}={(\omega_{\pm}-\omega_{\mathrm{ep}})}/{\omega_{\mathrm{ep}}}$ for varying $\gamma_{2}$ with $\gamma_{1}=-\gamma_{2}+\alpha\Gamma_{\mathrm{ep}}$ in a Drude sphere dimer with $k_{\mathrm{p}2}a=0.1$, $k_{\mathrm{p}1}a=k_{\mathrm{p}1,\mathrm{ep}}a$, and $d=5a$. Results for $\alpha\in\{0.95,1.00,1.05\}$ show eigenfrequency coalescence only at $\alpha=1$. The EPD is marked with $\circ$.
As discussed in Sec. III, $\mathscr{P}\mathscr{T}$-symmetry can be realized in the quasistatic limit by enforcing (17). For Drude dispersions, this corresponds to particles with identical plasma frequencies, $\omega_{\mathrm{p}1}=\omega_{\mathrm{p}2}=\omega_{\mathrm{p}}$, and damping rates of equal magnitude and opposite sign, $\gamma_{1}=-\gamma$ and $\gamma_{2}=\gamma$, with $\gamma>0$. In this configuration, the characteristic equation (14) reduces to
$\left(\frac{\omega}{\omega_{0}}\right)^{4}+\omega_{0}^{2}(R^{\mathrm{m}})^{2}C^{\mathrm{ed}}\left(\frac{\omega}{\omega_{0}}\right)^{2}+1-\left(\frac{C^{\mathrm{ed}}}{C_{M}}\right)^{2}=0,$ (22)
where $\omega_{0}=\omega_{\mathrm{p}}/\sqrt{3}$ is the isolated-sphere electric-dipole resonance frequency; $C^{\mathrm{ed}}$ and $C_{M}$ are the electric dipole and mutual capacitances, defined in Eqs. (47) and (50), respectively. For the longitudinal polarization, Eq. (22) is rewritten as
@@
where $\hat{\gamma}=\gamma/\omega_{0}$. Equation (23) is bi-quadratic in $(\omega/\omega_{0})$, hence eigenfrequencies appear in pairs $\pm(\omega/\omega_{0})$. Since the coefficients are real, any complex roots appear as conjugate pairs. The physical relevant roots with positive real part are
@@
In the absence of damping/gain ($\hat{\gamma}=0$), Eq. (25) is satisfied only by the trivial uncoupled limit $\rho=0$ consisting of widely separated particles that do not have an EPD. Alternatively, Eq. (25) is solved for $\rho$, providing the critical geometric condition necessary to realize an EPD for a given value of damping/gain.
@@
In the full-wave regime, the condition $\chi_{1}=\chi_{2}^{*}$ no longer guarantees a $\mathscr{P}\mathscr{T}$-symmetric configuration because of retardation and radiation losses, as discussed in Sec. III. In particular, if $\gamma$ is the only control parameter and all other parameters are fixed, the conjugate-symmetric choice $\chi_{1}=\chi_{2}^{*}$ does not produce an EPD: the eigenfrequency trajectories in the complex plane do not coalesce as $\gamma$ varies, as illustrated in Fig. 3(a).
@@
Figure 4: (a) Real and (b) imaginary parts of the normalized eigenfrequencies deviation $\delta\omega_{\pm}/\omega_{\mathrm{ep}}$ of the dimer ($k_{\mathrm{p}2}a=0.1$, $d=5a$) as a function of the normalized damping ratio detuning $\delta\gamma_{2}/\gamma_{2,\mathrm{ep}}=(\gamma_{2}-\gamma_{2,\mathrm{ep}})/\gamma_{2,\mathrm{ep}}$. The gain is tuned as $\gamma_{1}=-\gamma_{2}+\Gamma_{\mathrm{ep}}$. The EPD is marked with $\circ$.
@@
To synthesize an exceptional point in the full-wave regime, we fix $k_{\mathrm{p}2}=k_{\mathrm{p}}$ and $\rho=a/d$, and solve the system in Eq. (15), where the characteristic equation is given in Eq. (19). This yields the pair $(\omega_{\mathrm{ep}},\gamma_{2,\mathrm{ep}})$. The system is solved numerically by employing a root-finding algorithm, using the quasistatic solution $(\omega_{\mathrm{ep}}^{\mathrm{qs}},\gamma_{\mathrm{ep}}^{\mathrm{qs}})$ as the initial guess. Using Eq. (20), we then compute the effective resistance and inductance at the exceptional point, as $R_{1,\mathrm{ep}}^{\mathrm{m}}=\Re\!\left\{Z^{\mathrm{m}}_{1}(\omega_{\mathrm{ep}};\boldsymbol{\vartheta}_{\mathrm{ep}})\right\}$, and $L_{1,\mathrm{ep}}^{\mathrm{m}}={\Im\!\left\{Z^{\mathrm{m}}_{1}(\omega_{\mathrm{ep}};\boldsymbol{\vartheta}_{\mathrm{ep}})\right\}}/{\omega_{\mathrm{ep}}},$ which directly determine the Drude parameters of the gain particle as
$\omega_{\mathrm{p}1,\mathrm{ep}}=\frac{1}{\sqrt{L_{1,\mathrm{ep}}^{\mathrm{m}}\,\varepsilon_{0}a}},\qquad\gamma_{1,\mathrm{ep}}=\frac{R_{1,\mathrm{ep}}^{\mathrm{m}}}{L_{\mathrm{m},\mathrm{ep}}^{1}}.$ (27)
@@
It is important to emphasize that, unlike in the quasistatic limit, the quantity $\Gamma_{\mathrm{ep}}=\gamma_{1,\mathrm{ep}}+\gamma_{2,\mathrm{ep}}$ does not vanish. This construction guarantees the emergence of an exceptional point of degeneracy at $\omega=\omega_{\mathrm{ep}}$, as shown in Fig. 3(b). Here the normalized eigenfrequency deviation of the dimer ($k_{\mathrm{p}2}a=0.1$, $d=5a$) $\delta\omega_{\pm}/\omega_{\mathrm{ep}}=(\omega_{\pm}-\omega_{\mathrm{ep}})/\omega_{\mathrm{ep}}$ is plotted in the complex-frequency plane by varying $\gamma_{2}$ together with $\gamma_{1}$ enforcing the constraint $\gamma_{1}+\gamma_{2}=\alpha\Gamma_{\mathrm{ep}}$. The eigenfrequencies are obtained by numerically solving the characteristic equation (14). The plasma frequency of the gain particle is fixed at its EPD values, i.e. $k_{\mathrm{p}1}a=k_{\mathrm{p}1,\mathrm{ep}}a$. Results are shown for $\alpha\in\{0.95,1.00,1.05\}$. The plot highlights that the eigenfrequencies coalesce only for $\alpha=1$, where $\gamma_{1}=\gamma_{1,\mathrm{ep}}$ when $\gamma_{2}=\gamma_{2,\mathrm{ep}}$, while for $\alpha\neq 1$ the eigenfrequencies remain distinct and no coalescence is observed.
@@
For the case $\alpha=1$, Fig. 4 shows the normalized eigenfrequencies’ deviation $\delta\omega_{\pm}/\omega_{\mathrm{ep}}$ of the dimer as a function of the normalized damping ratio detuning $\delta\gamma_{2}/\gamma_{2,\mathrm{ep}}$, with the gain tuned according to $\gamma_{1}+\gamma_{2}=\Gamma_{\mathrm{ep}}$. The eigenfrequencies coalesce at $(\omega_{\mathrm{ep}},\gamma_{2,\mathrm{ep}})$ and exhibit the characteristic square-root splitting away from the EPD, a hallmark of exceptional point degeneracies.
@@ IV.2 Optical efficiencies in the proximity of EPDs
Figure 5: Peak’s frequency splitting $\Delta\omega_{\mathrm{pk}}/\omega_{0}$ versus normalized detuning $\delta d/d_{\mathrm{ep}}$ for the geometric control parameter $\vartheta_{i}=d/a$. Solid: CDA circuit model; dashed: multiparticle Mie theory. The value $d=0.99d_{\mathrm{ep}}$ (red $\times$), $d=d_{\mathrm{ep}}$ (blue $\times$), and $d=d_{\mathrm{ep}}$ (yellow $\times$) are provided for reference.
@@
Specifically, we quantify the sensitivity of the scattering efficiency to single-parameter perturbations in the vicinity of the EPD. Let the dimer be tuned to $(\omega_{\mathrm{ep}},\boldsymbol{\vartheta}_{\mathrm{ep}})$ and perturb a single control parameter, holding all others at their EPD values. As shown in Appendix H, the sensitivity of the system’s response, quantified here by the frequency splitting $\Delta\omega_{\mathrm{r}}:=\Re\{\omega_{+}-\omega_{-}\}$, diverges as the EPD is approached. The frequency-splitting behavior reveals two distinct regimes, depending on the sign of the perturbation $\delta\vartheta_{i}$. For negative perturbations $\delta\vartheta_{i}<0$, the system has two distinct resonance frequencies, and the splitting exhibits the characteristic square-root scaling $\Delta\omega_{\mathrm{r}}\sim|\delta\vartheta_{i}|^{1/2}$. Conversely, for $\delta\vartheta_{i}\geq 0$, the two resonance frequencies coalesce; in this regime, the difference manifests predominantly in the imaginary parts of the eigenfrequencies, associated with peak broadening, while the frequency splitting vanishes, i.e. $\Delta\omega_{\mathrm{r}}\approx 0$.
@@
This theoretical prediction is verified in Fig. 5. In this example, we consider a plasmonic dimer composed of two Drude spheres of radius $a$, separated by an inter-center distance $d=5a$. The system is illuminated by a plane wave with angular frequency $\omega$, polarized along the dimer axis and propagating along the direction perpendicular to the dimer plane. The lossy sphere is characterized by $k_{\mathrm{p}2}a=0.1$ and a damping rate $\gamma_{2}=\gamma_{2,\mathrm{ep}}$, while the gain sphere is tuned according to $\gamma_{1}=\gamma_{1,\mathrm{ep}}$ and $k_{\mathrm{p}1}=k_{\mathrm{p}1,\mathrm{ep}}$. Starting from the configuration $\boldsymbol{\vartheta}=\boldsymbol{\vartheta}_{\mathrm{ep}}$, we vary solely the interparticle distance (i.e. $\vartheta_{i}=d/a$) around the EPD value $d_{\mathrm{ep}}=5a$, and compute the peak-frequencies splitting $\Delta\omega_{\mathrm{pk}}=\omega_{\mathrm{pk}+}-\omega_{\mathrm{pk}-}$, where $\omega_{\mathrm{pk}+}$ and $\omega_{\mathrm{pk}-}$ are the high- and low-frequency peaks evaluated on the efficiencies spectra (e.g. Fig. 6), as a function of the normalized detuning $\delta d/d_{\mathrm{ep}}=(d-d_{\mathrm{ep}})/d_{\mathrm{ep}}$. We compare the predictions of the coupled-dipole model (solid line), with those obtained from the full wave multiparticle Mie Theory (dashed line), finding excellent agreement between the two approaches. Both numerical results clearly exhibit the predicted square-root behavior. This analysis confirms the characteristic square-root scaling also for the peak frequency splitting $\Delta\omega_{\mathrm{pk}}\sim|\delta d|^{1/2}$.
Additionally, for three representative values of $d/d_{\mathrm{ep}}\in\{0.99,1.00,1.01\}$ indicated by $\times$ in Fig. 5, Figure 6 shows the three efficiencies as functions of the normalized frequency detuning $\delta\omega/\omega_{\mathrm{ep}}={(\omega-\omega_{\mathrm{ep}})}/{\omega_{\mathrm{ep}}}$. The EPD transition is clear in Fig. 6. For $d=0.99\,d_{\mathrm{ep}}$ (red curve), two well-separated resonances appear, associated with eigenfrequencies with distinct real parts. At the critical value $d=d_{\mathrm{ep}}$ (blue curve), the two resonances coalesce into a single peak whose amplitude diverges, signaling eigenfrequencies coalescence. For $d=1.01\,d_{\mathrm{ep}}$ (yellow curve), the splitting is in the imaginary parts of the eigenfrequencies, and the response reduces to a single broadened resonance with reduced amplitude due to finite imaginary parts. In all three scenarios considered, $|Q_{\mathrm{e}}|$ is several orders of magnitude smaller than $Q_{\mathrm{s}}$ and $|Q_{\mathrm{a}}|$, due to the interplay between gain and loss. The results are in good agreement with the multiparticle Mie theory (dashed line), across all cases, confirming the accuracy of the proposed coupled-dipole circuit model.

### v2

@@ IV.1 Exceptional Points
Figure 3: (a) Normalized eigenfrequency deviation ${\delta\omega_{\mathrm{qs}}}/{\omega_{\mathrm{ep}}^{\mathrm{qs}}}={(\omega_{\pm}-\omega_{\mathrm{ep}}^{\mathrm{qs}})}/{\omega_{\mathrm{ep}}^{\mathrm{qs}}}$ in the complex-frequency plane for varying $\gamma_{2}=\gamma=-\gamma_{1}$, comparing full-wave (FW, blue) and quasistatic (QS, red) models for a conjugate-symmetric ($\chi_{1}=\chi_{2}^{*}$) dimer of Drude spheres with $k_{\mathrm{p}1}a=k_{\mathrm{p}2}a=0.1$ and $d=5a$. (b) Normalized eigenfrequency deviations in the FW regime ${\delta\omega}/{\omega_{\mathrm{ep}}}={(\omega_{\pm}-\omega_{\mathrm{ep}})}/{\omega_{\mathrm{ep}}}$ for varying $\gamma_{2}$ with $\gamma_{1}=-\gamma_{2}+\alpha\Gamma_{\mathrm{ep}}$ in a Drude sphere dimer with $k_{\mathrm{p}2}a=0.1$, $k_{\mathrm{p}1}a=k_{\mathrm{p}1,\mathrm{ep}}a$, and $d=5a$. Results for $\alpha\in\{0.95,1.00,1.05\}$ show eigenfrequency coalescence only at $\alpha=1$. The EPD is marked with $\circ$.
As discussed in Sec. III, $\mathscr{P}\mathscr{T}$-symmetry can be realized in the quasistatic limit by enforcing (17). For the Drude dispersions, this corresponds to particles with identical plasma frequencies, $\omega_{\mathrm{p}1}=\omega_{\mathrm{p}2}=\omega_{\mathrm{p}}$, and damping rates of equal magnitude and opposite sign, $\gamma_{1}=-\gamma$ and $\gamma_{2}=\gamma$, with $\gamma>0$. In this configuration, the characteristic equation (14) reduces to
$\left(\frac{\omega}{\omega_{0}}\right)^{4}+\left[(\omega_{0}R^{\mathrm{m}}C^{\mathrm{ed}})^{2}-2\right]\left(\frac{\omega}{\omega_{0}}\right)^{2}+1-\left(\frac{C^{\mathrm{ed}}}{C_{M}}\right)^{2}=0,$ (22)
where $\omega_{0}=\omega_{\mathrm{p}}/\sqrt{3}$ is the isolated-sphere electric-dipole resonance frequency; $C^{\mathrm{ed}}$ and $C_{M}$ are the electric-dipole and mutual capacitances, defined in Eqs. (47) and (50), respectively. For the longitudinal polarization, Eq. (22) is rewritten as
@@
where $\hat{\gamma}=\gamma/\omega_{0}$. Equation (23) is bi-quadratic in $(\omega/\omega_{0})$, hence eigenfrequencies appear in pairs $\pm(\omega/\omega_{0})$. Since the coefficients are real, any complex roots appear as conjugate pairs. The physically relevant roots with positive real part are
@@
In the absence of damping/gain ($\hat{\gamma}=0$), Eq. (25) is satisfied only by the trivial uncoupled limit $\rho=0$ consisting of widely separated particles that do not have an EPD. Alternatively, Eq. (25) is solved for $\rho$, providing the critical geometric condition necessary to realize an EPD for a given value of damping/gain. Notably, Eq. (25) admits two real solutions for $\hat{\gamma}^{2}$ for any physically admissible value of $\rho$. Indeed, the discriminant is $\Delta=16(1-4\rho^{6})$, which remains strictly positive under the non-overlapping condition $d>2a$ (i.e., $\rho<1/2$). In what follows, we focus on the minus-sign branch of Eq. (26), which corresponds to the solution with the smallest gain/loss level. In the full-wave regime, the condition $\chi_{1}=\chi_{2}^{*}$ no longer guarantees a $\mathscr{P}\mathscr{T}$-symmetric configuration because of retardation and radiation losses, as discussed in Sec. III. In particular, if $\gamma$ is the only control parameter and all other parameters are fixed, the conjugate-symmetric choice $\chi_{1}=\chi_{2}^{*}$ does not produce an EPD: the eigenfrequency trajectories in the complex plane do not coalesce as $\gamma$ varies, as illustrated in Fig. 3(a).
@@
Figure 4: (a) Real and (b) imaginary parts of the normalized eigenfrequency deviations $\delta\omega_{\pm}/\omega_{\mathrm{ep}}$ of the dimer ($k_{\mathrm{p}2}a=0.1$, $d=5a$) as a function of the normalized damping ratio detuning $\delta\gamma_{2}/\gamma_{2,\mathrm{ep}}=(\gamma_{2}-\gamma_{2,\mathrm{ep}})/\gamma_{2,\mathrm{ep}}$. The gain is tuned as $\gamma_{1}=-\gamma_{2}+\Gamma_{\mathrm{ep}}$. The EPD is marked with $\circ$.
@@
To synthesize an exceptional point in the full-wave regime, we fix $k_{\mathrm{p}2}=k_{\mathrm{p}}$ and $\rho=a/d$. Upon imposing Eq. (18), we determine $(\omega_{\mathrm{ep}},\gamma_{2,\mathrm{ep}})$ by requiring Eq. (19) and its derivative with respect to $\omega$ to vanish simultaneously. The system is solved numerically by employing a root-finding algorithm, using the quasistatic solution $(\omega_{\mathrm{ep}}^{\mathrm{qs}},\gamma_{\mathrm{ep}}^{\mathrm{qs}})$ as the initial guess. Using Eq. (20), we then compute the effective resistance and inductance at the exceptional point, as $R_{1,\mathrm{ep}}^{\mathrm{m}}=\operatorname{Re}\!\left\{Z^{\mathrm{m}}_{1}(\omega_{\mathrm{ep}};\boldsymbol{\vartheta}_{\mathrm{ep}})\right\}$, and $L_{1,\mathrm{ep}}^{\mathrm{m}}={\operatorname{Im}\!\left\{Z^{\mathrm{m}}_{1}(\omega_{\mathrm{ep}};\boldsymbol{\vartheta}_{\mathrm{ep}})\right\}}/{\omega_{\mathrm{ep}}},$ which directly determine the Drude parameters of the gain particle as
$\omega_{\mathrm{p}1,\mathrm{ep}}=\frac{1}{\sqrt{L_{1,\mathrm{ep}}^{\mathrm{m}}\,\varepsilon_{0}a}},\qquad\gamma_{1,\mathrm{ep}}=\frac{R_{1,\mathrm{ep}}^{\mathrm{m}}}{L_{1,\mathrm{ep}}^{\mathrm{m}}}.$ (27)
@@
It is important to emphasize that, unlike in the quasistatic limit, the quantity $\Gamma_{\mathrm{ep}}=\gamma_{1,\mathrm{ep}}+\gamma_{2,\mathrm{ep}}$ does not vanish. This construction yields an exceptional point of degeneracy at $\omega=\omega_{\mathrm{ep}}$, as shown in Fig. 3(b). Here, the normalized eigenfrequency deviation of the dimer ($k_{\mathrm{p}2}a=0.1$, $d=5a$) $\delta\omega_{\pm}/\omega_{\mathrm{ep}}=(\omega_{\pm}-\omega_{\mathrm{ep}})/\omega_{\mathrm{ep}}$ is plotted in the complex-frequency plane by varying $\gamma_{2}$ together with $\gamma_{1}$ enforcing the constraint $\gamma_{1}+\gamma_{2}=\alpha\Gamma_{\mathrm{ep}}$. The eigenfrequencies are obtained by numerically solving the characteristic equation (14). The plasma frequency of the gain particle is fixed at its EPD value, i.e. $k_{\mathrm{p}1}a=k_{\mathrm{p}1,\mathrm{ep}}a$. Results are shown for $\alpha\in\{0.95,1.00,1.05\}$. The plot highlights that the eigenfrequencies coalesce only for $\alpha=1$, where $\gamma_{1}=\gamma_{1,\mathrm{ep}}$ when $\gamma_{2}=\gamma_{2,\mathrm{ep}}$, while for $\alpha\neq 1$ the eigenfrequencies remain distinct and no coalescence is observed.
@@
For the case $\alpha=1$, Fig. 4 shows the normalized eigenfrequency deviation $\delta\omega_{\pm}/\omega_{\mathrm{ep}}$ of the dimer as a function of the normalized damping ratio detuning $\delta\gamma_{2}/\gamma_{2,\mathrm{ep}}$, with the gain tuned according to $\gamma_{1}+\gamma_{2}=\Gamma_{\mathrm{ep}}$. The eigenfrequencies coalesce at $(\omega_{\mathrm{ep}},\gamma_{2,\mathrm{ep}})$ and exhibit the characteristic square-root splitting away from the EPD, a hallmark of exceptional point degeneracies.
@@ IV.2 Optical efficiencies in the proximity of EPDs
Figure 5: Peak-frequency splitting $\Delta\omega_{\mathrm{pk}}/\omega_{0}$ versus normalized detuning $\delta d/d_{\mathrm{ep}}$ for the geometric control parameter $\vartheta_{i}=d/a$. Solid: CDA circuit model; dashed: multiparticle Mie theory. The values $d=0.99d_{\mathrm{ep}}$ (red $\times$), $d=d_{\mathrm{ep}}$ (blue $\times$), and $d=1.01d_{\mathrm{ep}}$ (yellow $\times$) are provided for reference.
@@
Specifically, we quantify the sensitivity of the scattering efficiency to single-parameter perturbations in the vicinity of the EPD. Let the dimer be tuned to $(\omega_{\mathrm{ep}},\boldsymbol{\vartheta}_{\mathrm{ep}})$ and perturb a single control parameter, holding all others at their EPD values. As shown in Appendix H, the sensitivity of the system’s response, quantified here by the frequency splitting $\Delta\omega_{\mathrm{r}}:=\operatorname{Re}\{\omega_{+}-\omega_{-}\}$, diverges as the EPD is approached. The frequency-splitting behavior reveals two distinct regimes, depending on the sign of the perturbation $\delta\vartheta_{i}$. For negative perturbations $\delta\vartheta_{i}<0$, the system has two distinct resonance frequencies, and the splitting exhibits the characteristic square-root scaling $\Delta\omega_{\mathrm{r}}\sim|\delta\vartheta_{i}|^{1/2}$. Conversely, for $\delta\vartheta_{i}\geq 0$, the two resonance frequencies coalesce; in this regime, the difference manifests predominantly in the imaginary parts of the eigenfrequencies, associated with peak broadening, while the frequency splitting vanishes, i.e. $\Delta\omega_{\mathrm{r}}\approx 0$.
This theoretical prediction is verified in Fig. 5. In this example, we consider a plasmonic dimer composed of two Drude spheres of radius $a$, separated by an inter-center distance $d=5a$. The system is illuminated by a plane wave with angular frequency $\omega$, polarized along the dimer axis and propagating along the direction perpendicular to the dimer plane. The lossy sphere is characterized by $k_{\mathrm{p}2}a=0.1$ and a damping rate $\gamma_{2}=\gamma_{2,\mathrm{ep}}$, while the gain sphere is tuned according to $\gamma_{1}=\gamma_{1,\mathrm{ep}}$ and $k_{\mathrm{p}1}=k_{\mathrm{p}1,\mathrm{ep}}$. Starting from the configuration $\boldsymbol{\vartheta}=\boldsymbol{\vartheta}_{\mathrm{ep}}$, we vary solely the interparticle distance (i.e. $\vartheta_{i}=d/a$) around the EPD value $d_{\mathrm{ep}}=5a$, and compute the peak-frequency splitting $\Delta\omega_{\mathrm{pk}}=\omega_{\mathrm{pk}+}-\omega_{\mathrm{pk}-}$, where $\omega_{\mathrm{pk}+}$ and $\omega_{\mathrm{pk}-}$ are the high- and low-frequency peaks evaluated on the efficiency spectra (e.g. Fig. 6), as a function of the normalized detuning $\delta d/d_{\mathrm{ep}}=(d-d_{\mathrm{ep}})/d_{\mathrm{ep}}$. We compare the predictions of the coupled-dipole circuit model (solid line), with those obtained from the full-wave multiparticle Mie theory (dashed line), finding excellent agreement between the two approaches. Both numerical results clearly exhibit the predicted square-root behavior. This analysis confirms the characteristic square-root scaling also for the peak-frequency splitting $\Delta\omega_{\mathrm{pk}}\sim|\delta d|^{1/2}$.
@@
Additionally, for three representative values of $d/d_{\mathrm{ep}}\in\{0.99,1.00,1.01\}$ indicated by $\times$ in Fig. 5, Figure 6 shows the three efficiencies as functions of the normalized frequency detuning $\delta\omega/\omega_{\mathrm{ep}}={(\omega-\omega_{\mathrm{ep}})}/{\omega_{\mathrm{ep}}}$. The EPD transition is clear in Fig. 6. For $d=0.99\,d_{\mathrm{ep}}$ (red curve), two well-separated resonances appear, associated with eigenfrequencies with distinct real parts. At the critical value $d=d_{\mathrm{ep}}$ (blue curve), the two resonances coalesce into a single peak with a strongly enhanced amplitude, signaling coalescence of the eigenfrequencies. For $d=1.01\,d_{\mathrm{ep}}$ (yellow curve), the splitting is in the imaginary parts of the eigenfrequencies, and the response reduces to a single broadened resonance with reduced amplitude due to finite imaginary parts. In all three scenarios considered, $|Q_{\mathrm{e}}|$ is several orders of magnitude smaller than $Q_{\mathrm{s}}$ and $|Q_{\mathrm{a}}|$, due to the interplay between gain and loss. The results are in good agreement with the multiparticle Mie theory (dashed line), across all cases, confirming the accuracy of the proposed coupled-dipole circuit model.